ya06120:

what is 5/7 divided by 35/8

Shadow:

Hello

ya06120:

hi?

Shadow:

\[\frac{ 5 }{ 7 } \div \frac{ 35 }{ 8 } = \frac{ 5 }{ 7 } \times \frac{ 8 }{ 35 }\]

Shadow:

You can do this one of two ways. Simplify first then multiply across. Or multiply across then simplify.

Shadow:

Five goes into 35 seven times, so we get: \[\frac{ 1 }{ 7 } \times \frac{ 8 }{ 7 }\]

ya06120:

kk

ya06120:

thanks for helping me

Shadow:

No problem. You are good from here right? Just have to multiply across (:

ya06120:

yeah ok ;-)